Accelerator Update for June 9, 2008 - June 11, 2008
Monday June 9
The day shift began with Operations monitoring Tevatron (TeV) store 6210 (85.54E30) and the Recycler (RR) stash (71.9E10), with the Antiproton Source (Pbar) stacking (34.6E10), with the Main Injector, Booster, and Linac operating normally, with the Collider Detector at Fermilab (CDF) and DZero (D0) using the colliding beams, and with the SciBooNE / MiniBooNE (BNB) experiments taking beam.
At 8:26 AM, a Linac RF station (LRF5) tripped off. It reset.
Operations put Pbar into standby at 8:42 AM for the duration of the Booster 95 LCW pump work.
The Booster 95 LCW pump switch was completed and Pbar resumed stacking at 9 AM.
At 10:35 AM, MIRF2 tripped off for the first of two times in nine minutes. It reset.
Operations began recovering the TeV at 12:02 PM.
Cryo system experts reported at 12:23 PM that the TeV A1 wet engine trip was a glitch. Cryo experts had the engine back on shortly after Operations terminated the store.
Operations prepared D0 for an access at 12:40 PM.
At 1:12 PM, a Linac Klystron RF station (KRF5) tripped off. It reset.
Operations put the TeV through a dry squeeze at 2:47 PM.
Operations reported at 4:01 PM that the Recycler had stashed antiprotons five times during the day shift.
At 4:09 PM, operators reported that they had lost the second transfer and half of the third transfer of antiprotons into the Recycler. Recycler experts investigated.
At 4:18 PM, Operations reported that the BNB horn tripped off. It reset.
At 5:23 PM, a Booster vacuum pump (IP132) tripped off. It reset.
At 5:52 PM, LRF1 tripped off. It reset.
Operations put the TeV into shot setup at 6:52 PM.
At 7:40 PM, a Recycler RF station (RRRF4) tripped off. It reset.
A Recycler expert began a RF study at 8:13 PM.
Operations established store 6213 at 8:21 PM with an initial luminosity of 252.47E30.
Operations reported at midnight that the Recycler had stashed four times during the evening shift.
Tuesday June 10
The midnight shift began with Operations monitoring store 6213 (133.93E30) and the stash (91.6E10), with Pbar stacking (40.3E10), and with BNB taking beam.
At 12:36 AM, Operations reported that a Recycler RF station (RRRF4) had tripped off several times. It reset.
At 12:55 AM, LRF4 tripped off. It reset
At 12:57 AM, KRF3 and 6 tripped off. They both reset.
At 4:52 AM, two Booster beam valves (VLNLAM and VLNDMP) closed. They both opened.
Operations reported at 7:58 AM that the Recycler had stashed five times during the midnight shift.
FESS personnel reported at 9:07 AM that they were beginning to clean all the MI air conditioning units and pond water strainers, starting with the MI-8 service building.
At 9:34 AM, a Switchyard (SY) vertical trim power supply (S:V204) tripped off. It reset. Operations contacted a power supply expert who investigated and reported that the supply had a plugged LCW line. He and another expert began the repair at 11:18 AM.
At 11:27 AM, KRF7 tripped off. It reset.
Operations terminated store 6213 at noon.
Power supply experts reported at 12:10 PM that they had completed work on S:V204.
Operations put the TeV into shot setup at 12:12 PM.
At 12:13 PM, Operations reported that the P1 transfer line beam permit had been tripping off due to the current of a correction magnet power supply (HT701). Operations restored previous files and adjusted a Lambertson (LAM52), but nothing would keep the trim in tolerance. Operations contacted an expert.
At 1:03 PM, a Pbar Debuncher kicker (D:IKIK) tripped off and would not reset. Operations contacted an expert.
Booster RF experts began bringing down the RF power supplies at 1:27 PM in preparation for CUB 95 LCW pump work and for heat exchanger work.
A Linac expert began work on KRF4 at 1:31 PM.
The Preaccelerator expert came out of the I- Source at 1:41 PM.
A Pbar expert reported at 1:42 PM that D:IKIK was back online. He found and replaced a bad power supply.
At 1:43 PM, Operations reported a large transverse oscillation in MI at injection. Experts found a bad control card for a power supply (I:H332).
MI experts began work on MP03 at 1:44 PM.
Operations had Booster prepared for access by 1:45 PM.
Operations had the MI-31 enclosure ready for access by 1:47 PM.
At 1:59 PM, RRRF4 tripped off. An expert replaced its power supply.
A Proton Source expert reported at 2:05 PM that he found a LCW leak on trim magnet (H803) in the MI-8 transfer line.
Operations established store 6214 at 2:09 PM with an initial luminosity of 213.3E30.
A RF expert reported at 2:19 PM that he found a LCW leak on BRF3. He estimated that it was losing about six gallons an hour. He began repairs.
Operations prepared MI for access at 2:23 PM.
At 2:26 PM, D:IKIK stopped pulsing.
At 2:34 PM, CDF reported a Tevatron Electron Lens (TEL) alarm. Operators found that it had tripped off along with a couple of corrector power supplies. All these devices had tripped off on a water flow indication. Operators cycled the LCW supply valve and everything reset.
A kicker expert reported at 3 PM that he had D:IKIK pulsing again.
FESS reported at 3:39 PM that they had completed their MI work.
Operations reported at 3:56 AM that the Recycler had stashed four times during the day shift.
A Booster expert reported at 4:03 PM that he had completed his work on MP03.
Linac experts reported at 4:06 PM that they had completed their work on KRF4 and LRF2 was being pumped down to its normal vacuum level.
Operations reported at 4:42 PM that Booster was ready for beam.
Pbar resumed stacking at 5:33 PM.
At 5:41 PM, a Booster vertical bend magnet power supply (B:VBC1) tripped off on overcurrent and would not reset. Booster experts were called in. They changed out a control module and a bad fan.
Operations resumed sending beam to Pbar, BNB, and MTest around 7:56 PM.
TeV experts began a one-hour study period at 8:55 PM.
Operations disabled the MTest beam at 8:58 PM.
At 10:10 PM, the P1 transfer line tripped off for the first of two times in two minutes due to a quadrupole magnet power supply (I:Q702) trip. Everything reset.
Operations reported at midnight that the Recycler had stashed three times during the evening shift.
Wednesday June 11
The midnight shift began with Operations monitoring store 6214 (75.67E30) and the stash (98.3E10), with Pbar stacking (33.0E10), and with BNB taking beam.
At 1:19 AM, KRF3 tripped off. It reset.
At 3:46 AM, KRF3 tripped off. It reset.
At 4:58 AM, Operations reported problems with the MI ramp. Operations restarted the ramp so it would learn in correctly.
At 5:35 AM, Operations reported a problem with the Pbar target. Operators halted stacking and contacted an expert.
Operations prepared Pbar for access at 7:17 AM.
At 7:26 AM, a Booster vacuum pump (IPRF17) tripped off and caused BRF7 & 8 to trip. Everything reset.
Operations reported at 7:57 AM that the Recycler had stashed five times during the midnight shift.
The Plans for Wednesday
The plan for today is to find and fix the Pbar target problem. This work might take eight to ten hours.
Comments and Suggestions
|last modified 06/11/2008 email Fermilab|
|Security, Privacy, Legal|